OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24. 18 32 64

OpenStudy (anonymous):

y = cx, where c is a constant 12 = c16 so c = 12/16 = 3/4 y=cx becomes y = 3/4 x if y = 24, then we have 24 = 3/4x x= 4/3(24) = 32